Accelerator Update for March 17, 2008 - March 19, 2008
Monday March 17
The day shift began with Operations monitoring the Tevatron (TeV) store 5989 (212.1E30) and the Recycler stash (113.2E10), with the Antiproton Source stacking (20.2E10), with the Main Injector (MI), Linac, and Booster being prepared for access, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with NuMI and the SciBooNE / MiniBooNE (BNB) experiments off due to accesses.
At 9:13 AM, a LCW pump at MI-20 tripped off. This was due to the switch to recirculation mode in MI. A water expert said the pump should be turned back on when the Recycler was empty of particles.
At 9:20 AM, CUB experts reported that the Booster 95 LCW pump (LP1) was failing.
Operations reported at 9:44 AM that CUB experts began the switch to the LP2 pump. Unfortunately, this switch did not go well. By the time they had finished, the Booster East RF system and the Tevatron Electron Lens (TEL1) had both tripped off due to a large pressure drop. Operations reported problems with bringing BRF2 & 6 back online.
Booster came back online at 10:11 AM.
At 10:22 AM, BRF11 tripped off. It reset.
At 10:49 AM, Operations reported that the MI water pressure had been dropping ever since the LCW system was put in recirculating mode. After a bit of investigation, Operations discovered that the water experts had made up for lost LCW.
A Pbar expert reported at 12:53 PM that the Sweeper “A” magnet was back up and running at AP0.
A Recycler expert reported at 12:58 PM that he had adjusted the skew quadrupole magnets in the 404 area.
Operations reported at 4:10 PM that the Recycler had stashed antiprotons four times during the day shift.
At 4:29 PM, BRF1 tripped off. It reset.
At 4:31 PM, BRF11 tripped off. It reset.
At 4:36 PM, BRF1 & 11 tripped off. They reset.
A MI expert began checking coalescing at 4:50 PM.
At 5:03 PM, BRF4 & 9 tripped off. They reset.
At 5:05 PM, BRF9 tripped off. It reset.
At 5:21 PM, BRF16 tripped off. It reset.
At 5:51 PM, Operations lowered the beam intensity for BNB due to Booster beam position problems. Operators investigated.
Linac resumed sending beam at 11:34 PM. Operations reported that the Pbar lithium lens started tripping off when Pbar tried to resume stacking. After several resets, the lens finally came on and stayed on.
Operations reported at 11:57 PM that the Recycler had stashed four times during the evening shift.
Tuesday March 18
The midnight shift began with Operations monitoring store 5989 (66.92E30) and the stash (343.8E10), with Pbar stacking (28.5E10), and with NuMI and BNB taking beam.
At 12:24 AM, BRF16 tripped off. It reset.
At 12:28 AM, a Booster power supply (B:CEP400) tripped off. It reset.
Operations terminated store 5989 at 3:13 AM.
Operations put the TeV into shot setup at 3:24 AM.
At 3:50 AM, KRF3 tripped off. It reset.
A TeV expert conducted an energy match between the TeV and MI at 4:14 AM.
Operations established store 5990 at 5:14 AM with an initial luminosity of 259.78E30.
At 5:23 AM, Operations reported that the TeV sector A17 separator sparked and caused a luminosity loss of about 50E30.
At 6:37 AM, LRF3 tripped off. It reset, but tripped off again. Operators thought they heard the station spark. Operations called in an expert.
At 7:05 AM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to trip also. Everything reset.
Operations reported at 7:36 AM that they turned on the MI-20 LCW pump #2.
At 7:37 AM, Operations reported a MI leak alarm. The water experts will make up water for MI.
Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.
A Linac expert reported at 8:22 AM that LRF3 was back online.
Operations had reestablished beam to Booster, Pbar, NuMI, and BNB by 8:26 AM.
Switchyard (SY) experts began sending beam to SY and Meson at 8:37 AM for MTest beam studies.
A controls expert reported at 9:41 AM that he discovered that the MI BPM problems were due to a “flakey” VME crate. It will take him an hour to replace the crate. The MI-8 autotune will be down during this work.
Recycler experts completed a damper study at 10:42 AM.
At 10:59 AM, Operations halted the MI ramp to investigate the MI-30 quadrupole power supply for a possible LCW leak. The power supply expert reported that there was a small leak, but it wasn’t dripping on anything. He will do the repair tomorrow during the access.
MI resumed sending beam at 11:33 AM.
At 12:37 PM, KRF6 tripped off. It reset.
At 1 PM, the Linac Klystron vernier tripped off. It reset.
At 2:01 PM, Operations reported high temperatures for the MI-10 LCW system. An operator contacted experts who placed the MI-10 pond water pump into its high mode. The expert asked Operations to turn the pump back to low after the temperature drops.
At 2:30 PM, MIRF5 tripped off. It reset.
At 2:47 PM, Operations reported that the high voltage for the I- Source was alternating between two values. Operators changed its oscillator and control modules, which fixed the problem.
Operations reported at 4:02 PM that the Recycler had stashed four times during the day shift.
At 4:09 PM, Operations reported that vacuum problems in the Recycler were causing antiprotons to be lost in the 800 area.
Operations reported at 5:07 PM that they had switched the MI-50 pond water pump to low.
SY experts reported at 5:57 PM that they had completed their MTest studies for the day.
At 8:33 PM, BRF16 tripped off. It reset.
At 11:52 PM, a Booster vacuum pump (IPBD2) tripped off and caused two vacuum valves (VLNLAM and VLNDMP) to close. Everything reset and opened.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Wednesday March 19
The midnight shift began with Operations monitoring store 5990 (57.91E30) and the stash (145.9E10), with Pbar stacking (37.6E10), and with NuMI and BNB taking beam.
At 4:32 AM, LRF4 tripped off for the first of three times in about 90 minutes. It reset.
At 5:08 AM, KRF2 tripped off. It reset.
At 6:31 AM, LRF3 tripped off for the first of four times in about an hour due to line voltage irregularities. Operations reported that this same thing had happened yesterday at the same time.
Operations reported at 7:58 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store. The Run Coordinator will allow TeV and MI experts to conduct studies. The access planned for today was postponed to Thursday, 3/20/08.
Comments and Suggestions
|last modified 03/19/2008 email Fermilab|
|Security, Privacy, Legal|